# DCA: Graph-Guided Deep Embedding Clustering for Brain Atlases

### v1

@@ 5 Experimental results
To rigorously benchmark our method, we developed a comprehensive evaluation framework that assesses parcellation quality via intra- and inter-region fMRI signal correlations. Next, we apply each group atlas to a suite of real-world downstream tasks, such as cognitive task decoding to evaluate practical utility. We compare DCA against several widely used atlases: Yeo et al. yeo2011organization parcellates cortex into seven large-scale functional networks derived from resting-state connectivity. Brodmann brodmann1909vergleichende defined cortical areas based on cytoarchitectonic boundaries. Schaefer et al. schaefer2018local published a multi-resolution functional atlas (100–1000 parcels) using gradient-informed clustering of functional connectivity. AAL (Automated Anatomical Labeling) tzourio2002automated ; rolls2020automated and MUSE (MUlti-atlas region Segmentation utilizing Ensembles) doshi2016muse divide the brain into hundreds of anatomy-based regions. MMP (Multi-Modal Parcellation) glasser2016multi integrates structural, functional, and connectivity data to define 360 cortical areas. GIANT (Genetically Informed brAiN aTlas) bao2025genetically achieves brain parcellation by genetically-driven integration of voxel-wise heritability and spatial proximity. Additionally, the Watershed Atlas and the Allen Human Reference Atlas-3D, 2020, were incorporated. To study the effect of granularity, we further generate DCA group atlases with $K\in\{41,100,200,360,400,500,800\}$ parcels constrained to the FreeSurfer gray-matter mask. The mask is derived from the cortical gray matter regions in FreeSurfer’s aparc+aseg.mgz fischl2012freesurfer and transformed into MNI152 space. In the following sections, we report correlation metrics and downstream task accuracies for each atlas configuration.
@@ 5.1 Main results
Fig. 3 summarizes clustering performance for several existing atlases alongside our DCA method, evaluated by Homogeneity (Fig. 3B) and Silhouette Coefficient (Fig. 3C). We sample 100 subjects from HCP randomly. Full quantitative results are given in Table 4. As the number of parcels increases from 7 to 1000, both metrics rise for all methods. To ensure a fair comparison, we match each atlas to DCA at the similar ROI numbers. Across every evaluated resolution, DCA consistently outperforms the best-performing baseline atlas, yielding higher homogeneity and silhouette scores at each scale. For example, at 200 parcels, DCA improves Homogeneity by 77.7% and Silhouette by 19.5% over the Schaefer baseline. On average, across 41 to 800 parcels, DCA improves Homogeneity by 98.8% and Silhouette by 29%, demonstrating consistent gains in functional homogeneity and spatial separation. This highlights that DCA provides a more refined and effective parcellation solution across various atlas configurations, regardless of the cluster count.
@@ 5.2 Downstream tasks
We evaluate the utility of DCA atlases on six downstream classification tasks covering behavioral prediction, cognitive decoding, and clinical diagnosis on the same 100 subjects of previous section (Fig. 4). Across resolutions, DCA consistently matches or outperforms the strongest baseline within each group. Full results, including extended benchmarks on additional tasks, are reported in the Appendix (Table 8).
@@ 5.3 Task-specific atlas
We demonstrate that our framework can be readily adapted to task-specific settings, achieving substantial improvements on the corresponding evaluation metrics. Specifically, we replace the reconstruction self-supervised Swin-UNETR with a version fine-tuned for gender classification and use its encoder to derive a task-specific atlas, denoted $\mathrm{DCA}^{\text{gender}}_{100}$. We then aggregate fMRI signals into $K=100$ ROIs for $N=200$ subjects drawn from the Swin UNETR fine-tuning test split to avoid data leakage and evaluated using two downstream classifiers:
@@ 5.4 Ablation study
To quantify the contribution of each component, we conducted two ablation studies on on 100 random subjects from HCP. First, we applied (1) K-Means clustering and (2) the same graph construction and graph-cut pipeline from our method directly to the raw fMRI time series. This baseline highlights the necessity of our pretrained encoder for extracting informative features. Second, we ran both (3) K-Means and (4) the same graph cut pipeline on the encoded embeddings without the KL-guided joint optimization of encoder parameters and centroids, isolating the impact of our graph-guided deep clustering mechanism. In both ablations, all similarity metrics (homogeneity and silhouette) drop noticeably below those of (5) the full DCA model (Fig. 5 and Table 3). Moreover, K-Means fails to produce spatially contiguous parcels. And graph-cut improves continuity, but it still yields isolated regions. In contrast, DCA’s iterative KL-driven refinement—which alternates updating graph weights, encoder parameters, and cluster centers—produces brain atlases that are highly homogeneous. We also tested the model reproducibility (Table 6 and 7) , the effects of the loss (Table 15 and 20, Fig. 12) and normalization (Table 15), the smoothing (Table 16), graph cut method (Fig. 10 and Table 17), gray matter region (Table 18), number of neighbors (Table 19) and centroid initialization (Table 21) as detailed in Appendix.
@@
Table 3: Evaluation for fMRI and embedding-based clustering with KMeans, Graph Cut, and DCA.
@@ 6 Discussion & Conclusion
We further examined how downstream task performance varies with the number of parcels. Based on these trends, tasks can be broadly grouped into three categories: (i) resolution-optimal tasks, which peak at intermediate resolutions, (ii) resolution-insensitive tasks, which remain largely stable across scales, and (iii) size-driven tasks, which monotonically track parcel granularity (Table 13). Importantly, this pattern is consistently observed for both DCA and Schaefer atlases (Table 14), indicating that there is no universally optimal resolution. Instead, the choice of parcel count should be guided by the target application. In addition, DCA offers the flexibility to further optimize atlases for specific tasks at a given resolution, as demonstrated in our task-adapted atlas experiments (Table 1 and 2).
@@
Despite its strengths, DCA has several limitations. First, voxel-wise representation learning and clustering incur substantial memory and computational costs, especially at whole-brain scales. Future work may explore region-specific parcellation or sparse embedding schemes to reduce overhead. Second, our reliance on fixed KNN graphs to enforce spatial continuity may inadvertently attenuate long-range functional relationships, which may suppress long-range functional relationships. Integrating adaptive or learned graphs could help balance spatial continuity with network-level functional coherence. Lastly, the current pipeline uses only single-modality fMRI data. Incorporating structural and diffusion imaging, or even electrophysiological data (e.g. sEEG), could further enhance the biological fidelity of parcellations zhi2022evaluating ; glasser2016multi .Meanwhile, traditional atlas construction methods struggle to reconcile conflicting signals across modalities yan2023homotopic , motivating the development of a multimodal deep-clustering framework as a promising avenue for building richer, functionally and structurally grounded brain atlases.
@@ 8 Cross-dataset generalization on CHCP
We further assess our model’s cross-dataset generalization by applying the Swin-UNETR encoder—pretrained on the HCP dataset—to individual atlas generation on the other independent dataset (Fig. 6 and Table 5). The Chinese Human Connectome Project (CHCP) dataset comprises high-resolution multimodal MRI—including structural, diffusion, and resting-state fMRI—from healthy Chinese adults, and uses the same acquisition parameters and HCP preprocessing pipeline [52]. Our results demonstrate that, without any additional fine-tuning, DCA produces coherent, spatially contiguous parcellations on CHCP dataset that surpass other atlases in both metric, underscoring the robustness of the learned voxel embeddings.
@@ 10 Model reproducibility
While method produces consistent atlases when run multiple times on the same fMRI segment with identical settings, we then evaluated cross-segment stability. To quantify consistency, we produced $\mathrm{DCA}_{100}$ atlases from ten non-overlapping fMRI segments for each of ten HCP participants, spanning multiple runs and both phase-encoding directions. Atlas similarity was measured with Dice and intersection-over-union (IoU). As summarised in Table 7, intra-subject similarity is markedly higher than both inter-subject similarity and a null baseline obtained by randomly partitioning the cortical mask into 100 spatially contiguous, equal-sized parcels (with parcel correspondence solved via the Hungarian algorithm). These results confirm that DCA yields reproducible atlases despite stochastic initialisation.
@@ 11 Group atlas generation
12:  Keep largest 6-connected component of label $k$
@@ 12 AtlaScore: atlas evaluation platform
In addition to the 12 downstream tasks provided by AtlaScore, we further evaluated atlas utility using modern neural classifiers. Specifically, we trained a compact 1-D CNN on ROI-level time series (Tables LABEL:tab:lowLABEL:tab:high and Table 1) and a graph-based $k$-GNN on FC graphs (Table 2). These models provide a complementary perspective by directly testing whether atlas parcellations facilitate feature extraction for nonlinear learning. These findings reinforce that DCA not only benefits classical SVC pipelines but also enhances performance in deep learning–based settings.
@@
1 Similarity–Homogeneity Measure mean intra-cluster correlation
@@ 12.1 Distance-controlled boundary coefficient (DCBC)
By controlling for spatial distance in this way, DCBC disentangles true functional boundaries from artifacts of spatial smoothness, providing a more reliable parcellation assessment. In our evaluation process, all parameters followed the settings specified in [50]. However, we note that DCBC was mainly developed for surface-based parcellations and becomes computationally prohibitive at the fine voxel resolution employed by DCA. Therefore, we computed DCBC scores by projecting volumetric atlases onto the cortical surface (fsLR 32k template) [53], analyzing only data from the left hemisphere. This surface-based approach exceeds the scope of our native volumetric framework.
@@ 12.2 Gender classification
To ensure subject-level generalization, we performed 10-fold cross-validation across subjects: in each fold, 90 subjects were used for training and 10 for testing. A linear support vector classifier (SVC) was trained on the training set. If the FC dimensionality exceeded 100, we applied principal component analysis (PCA) to reduce dimensionality: features were projected onto the top 100 principal components computed from the training data, and test samples were projected into the same subspace. Classification accuracy on the test subjects was recorded for each fold and averaged to obtain final performance.
@@ 13 Complete DCA performance on AtlaScore
13.1 Evaluation on different smoothing level
@@
Figure 8: Homogeneity and silhouette measured over 100 HCP subjects on different smoothing level.
@@ 13.2 DCBC performance across Atlases
DCBC was mainly developed for surface-based parcellations and becomes computationally prohibitive at the fine voxel resolution employed by DCA. Therefore, we computed DCBC scores by projecting volumetric atlases onto the cortical surface (fsLR 32k template [53]), analyzing only data from the left hemisphere. This surface-based approach exceeds the scope of our native volumetric framework (Fig. 9 and Table 11 ).
@@
Atlas #Parcels mean st.d.
@@ 14 Effect of parcel number on downstream tasks
To examine whether downstream tasks prefer specific spatial resolutions, we varied the number of parcels for both DCA (41, 100, 200, 360, 400, 500) and Schaefer (100, 200, 300, 400, 500) atlases. For each resolution we performed subject-level cross-validation and report accuracies in Tables 13 and 14. Across both atlas families, three consistent patterns emerge:
@@
  • Resolution-insensitive tasks. Several clinical endpoints (e.g., AD diagnosis) exhibit weak or no monotonic trend across resolutions for both DCA and Schaefer. These tasks likely rely on subcortical or global signals that cortex-only atlases do not explicitly model, so varying cortical granularity alone has limited effect. In such settings, choosing resolution can be guided by computational cost or downstream interpretability rather than accuracy.
@@ 15.1 Regularization and reconstruction loss
We investigated whether incorporating an orthogonality regularizer or a reconstruction loss would further improve parcellation quality (Table 15). Applying these augmented objectives to the same 100 subjects used in our main experiments yielded no statistically significant gains in homogeneity or silhouette coefficient for 100 parcels, indicating that the core KL-based clustering loss is sufficient to drive optimal voxel-level atlas generation.
@@
The orthogonality regularizer is defined on the centroid matrix $\mathbf{D}\in\mathbb{R}^{K\times d}$ (with unit-norm rows) as follows. Let
@@
be the Gram matrix of pairwise inner products. We zero out the diagonal to isolate off-diagonal similarities:
@@
The orthogonality loss then penalizes the mean absolute off-diagonal entry via
@@ 15.3 Choice of graph cut method
In addition to spectral clustering, we evaluated several graph-cut algorithms (Fig. 10 and Table 17). However, most failed to guarantee that each resulting parcel forms a single connected subgraph, leading to fragmented regions. Here, we use a breadth-first search (BFS) based algorithm. The weighted BFS–connected clustering algorithm begins by converting the input edge list and weights into an undirected adjacency list, then randomly seeds $k$ initial clusters by assigning one unique node to each cluster. Each cluster maintains a max-heap of its unassigned neighboring nodes, prioritized by edge weight. Clusters then grow in parallel: at each step, a cluster pops the highest-weight neighbor from its heap, claims that node (if unassigned), and pushes all of its unassigned neighbors onto the heap. To enforce roughly equal cluster sizes, each cluster stops growing once it reaches $\lceil N/k\rceil$. If any nodes remain unassigned after this frontier-driven expansion, they are absorbed into the smallest adjacent cluster. By always selecting the strongest edges first and only adding connected nodes, this method produces contiguous clusters that respect the underlying graph’s weighted connectivity.

### v2

@@ 5 Experimental results
To rigorously benchmark our method, we developed a comprehensive evaluation framework that assesses parcellation quality via intra‐ and inter‐region fMRI signal correlations. Next, we apply each group atlas to a suite of real‐world downstream tasks, such as cognitive task decoding to evaluate practical utility. We compare DCA against several widely used atlases: Yeo et al. yeo2011organization parcellates cortex into seven large‐scale functional networks derived from resting‐state connectivity. Brodmann brodmann1909vergleichende defined cortical areas based on cytoarchitectonic boundaries. Schaefer et al. schaefer2018local published a multi‐resolution functional atlas (100–1000 parcels) using gradient‐informed clustering of functional connectivity. AAL (Automated Anatomical Labeling) tzourio2002automated ; rolls2020automated and MUSE (MUlti-atlas region Segmentation utilizing Ensembles) doshi2016muse divide the brain into hundreds of anatomy‐based regions. MMP (Multi‐Modal Parcellation) glasser2016multi integrates structural, functional, and connectivity data to define 360 cortical areas. GIANT (Genetically Informed brAiN aTlas) bao2025genetically achieves brain parcellation by genetically-driven integration of voxel-wise heritability and spatial proximity. Additionally, the Watershed Atlas and the Allen Human Reference Atlas-3D, 2020, were incorporated. To study the effect of granularity, we further generate DCA group atlases with $K\in\{41,100,200,360,400,500,800\}$ parcels constrained to the FreeSurfer gray‐matter mask. The mask is derived from the cortical gray matter regions in FreeSurfer’s aparc+aseg.mgz fischl2012freesurfer and transformed into MNI152 space. In the following sections, we report correlation metrics and downstream task accuracies for each atlas configuration.
@@ 5.1 Main results
Fig. 3 summarizes clustering performance for several existing atlases alongside our DCA method, evaluated by Homogeneity (Fig. 3B) and Silhouette Coefficient (Fig. 3C). We sample 100 subjects from HCP randomly. Full quantitative results are given in Table 4. As the number of parcels increases from 7 to 1000, both metrics rise for all methods. To ensure a fair comparison, we match each atlas to DCA at the similar ROI numbers. Across every evaluated resolution, DCA consistently outperforms the best‐performing baseline atlas, yielding higher homogeneity and silhouette scores at each scale. For example, at 200 parcels, DCA improves Homogeneity by 77.7% and Silhouette by 19.5% over the Schaefer baseline. On average, across 41 to 800 parcels, DCA improves Homogeneity by 98.8% and Silhouette by 29%, demonstrating consistent gains in functional homogeneity and spatial separation. This highlights that DCA provides a more refined and effective parcellation solution across various atlas configurations, regardless of the cluster count.
@@ 5.2 Downstream tasks
We evaluate the utility of DCA atlases on six downstream classification tasks covering behavioral prediction, cognitive decoding, and clinical diagnosis on the same 100 subjects as in the previous section (Fig. 4). Across resolutions, DCA consistently matches or outperforms the strongest baseline within each group. Full results, including extended benchmarks on additional tasks, are reported in the Appendix (Table 8).
@@ 5.3 Task-specific atlas
We demonstrate that our framework can be readily adapted to task-specific settings, achieving substantial improvements on the corresponding evaluation metrics. Specifically, we replace the reconstruction self-supervised Swin-UNETR with a version fine-tuned for gender classification and use its encoder to derive a task-specific atlas, denoted $\mathrm{DCA}^{\text{gender}}_{100}$. We then aggregate fMRI signals into $K=100$ ROIs for $N=200$ subjects drawn from the Swin-UNETR fine-tuning test split to avoid data leakage and evaluated using two downstream classifiers:
@@ 5.4 Ablation study
To quantify the contribution of each component, we conducted two ablation studies on 100 random subjects from HCP. First, we applied (1) K-Means clustering and (2) the same graph construction and graph-cut pipeline from our method directly to the raw fMRI time series. This baseline highlights the necessity of our pretrained encoder for extracting informative features. Second, we ran both (3) K-Means and (4) the same graph cut pipeline on the encoded embeddings without the KL-guided joint optimization of encoder parameters and centroids, isolating the impact of our graph-guided deep clustering mechanism. In both ablations, all similarity metrics (homogeneity and silhouette) drop noticeably below those of (5) the full DCA model (Fig. 5 and Table 3). Moreover, K-Means fails to produce spatially contiguous parcels. And graph-cut improves continuity, but it still yields isolated regions. In contrast, DCA’s iterative KL-driven refinement—which alternates updating graph weights, encoder parameters, and cluster centers—produces brain atlases that are highly homogeneous. We also tested the model reproducibility (Table 6 and 7), the effects of the loss (Table 15 and 20, Fig. 12) and normalization (Table 15), the smoothing (Table 16), graph cut method (Fig. 10 and Table 17), gray matter region (Table 18), number of neighbors (Table 19) and centroid initialization (Table 21) as detailed in Appendix.
@@
Table 3: Evaluation for fMRI and embedding-based clustering with K-Means, Graph Cut, and DCA.
@@ 6 Discussion & Conclusion
We further examined how downstream task performance varies with the number of parcels. Based on these trends, tasks can be broadly grouped into three categories: (i) resolution-optimal tasks, which peak at intermediate resolutions, (ii) resolution-insensitive tasks, which remain largely stable across scales, and (iii) size-driven tasks, which monotonically track parcel granularity (Table 13). Importantly, this pattern is consistently observed for both DCA and Schaefer atlases (Table 14), indicating that there is no universally optimal resolution. Instead, the choice of parcel count should be guided by the target application. In addition, DCA offers the flexibility to further optimize atlases for specific tasks at a given resolution, as demonstrated in our task-adapted atlas experiments (Table 1 and 2). What’s more, no single atlas generalizes optimally across tasks—different tasks favor different parcellations (salehi2020there, )—so developing task-specific atlases is warranted.
@@
Despite its strengths, DCA has several limitations. First, voxel-wise representation learning and clustering incur substantial memory and computational costs, especially at whole-brain scales. Future work may explore region-specific parcellation or sparse embedding schemes to reduce overhead. Second, our reliance on fixed KNN graphs to enforce spatial continuity may inadvertently attenuate long‐range functional relationships, which may suppress long-range functional relationships. Integrating adaptive or learned graphs could help balance spatial continuity with network-level functional coherence. Lastly, the current pipeline uses only single-modality fMRI data. Incorporating structural and diffusion imaging, or even electrophysiological data (e.g. sEEG), could further enhance the biological fidelity of parcellations zhi2022evaluating ; glasser2016multi .Meanwhile, traditional atlas construction methods struggle to reconcile conflicting signals across modalities yan2023homotopic , motivating the development of a multimodal deep‐clustering framework as a promising avenue for building richer, functionally and structurally grounded brain atlases.
@@ 8 Cross-dataset generalization on CHCP
We further assess our model’s cross‐dataset generalization by applying the Swin-UNETR encoder—pretrained on the HCP dataset—to individual atlas generation on the other independent dataset (Fig. 6 and Table 5). The Chinese Human Connectome Project (CHCP) dataset comprises high‐resolution multimodal MRI—including structural, diffusion, and resting‐state fMRI—from healthy Chinese adults, and uses the same acquisition parameters and HCP preprocessing pipeline [52]. Our results demonstrate that, without any additional fine‐tuning, DCA produces coherent, spatially contiguous parcellations on CHCP dataset that surpass other atlases in both metrics, underscoring the robustness of the learned voxel embeddings.
@@ 10 Model reproducibility
While the method produces consistent atlases when run multiple times on the same fMRI segment with identical settings, we then evaluated cross-segment stability. To quantify consistency, we produced $\mathrm{DCA}_{100}$ atlases from ten non-overlapping fMRI segments for each of ten HCP participants, spanning multiple runs and both phase-encoding directions. Atlas similarity was measured with Dice and intersection-over-union (IoU). As summarised in Table 7, intra-subject similarity is markedly higher than both inter-subject similarity and a null baseline obtained by randomly partitioning the cortical mask into 100 spatially contiguous, equal-sized parcels (with parcel correspondence solved via the Hungarian algorithm). These results confirm that DCA yields reproducible atlases despite stochastic initialisation.
@@ 11 Group atlas generation
12:  Keep the largest 6-connected component of label $k$
@@ 12 AtlaScore: atlas evaluation platform
In addition to the 12 downstream tasks provided by AtlaScore, we further evaluated atlas utility using modern neural classifiers. Specifically, we trained a compact 1-D CNN on ROI-level time series (Table 1) and a graph-based $k$-GNN on FC graphs (Table 2). These models provide a complementary perspective by directly testing whether atlas parcellations facilitate feature extraction for nonlinear learning. These findings reinforce that DCA not only benefits classical SVC pipelines but also enhances performance in deep learning–based settings.
@@
1 Similarity–Homogeneity Measure mean intra‐cluster correlation
@@ 12.1 Distance-controlled boundary coefficient (DCBC)
By controlling for spatial distance in this way, DCBC disentangles true functional boundaries from artifacts of spatial smoothness, providing a more reliable parcellation assessment. In our evaluation process, all parameters followed the settings specified in [50]. However, we note that DCBC was mainly developed for surface‐based parcellations and becomes computationally prohibitive at the fine voxel resolution employed by DCA. Therefore, we computed DCBC scores by projecting volumetric atlases onto the cortical surface (fsLR 32k template) [53], analyzing only data from the left hemisphere. This surface-based approach exceeds the scope of our native volumetric framework.
@@ 12.2 Gender classification
To ensure subject-level generalization, we performed 10-fold cross-validation across subjects: in each fold, 90 subjects were used for training and 10 for testing. A linear support vector classifier (SVC) was trained on the training set. If the FC dimensionality exceeded 100, we applied principal component analysis (PCA) to reduce dimensionality: features were projected onto the top 100 principal components computed from the training data, and test samples were projected into the same subspace. Classification accuracy on the test subjects was recorded for each fold and averaged to obtain the final performance.
@@ 13 Complete DCA performance on AtlaScore
13.1 Evaluation on different smoothing levels
@@
Figure 8: Homogeneity and silhouette were measured over 100 HCP subjects on different smoothing levels.
@@ 13.2 DCBC performance across Atlases
DCBC was mainly developed for surface‐based parcellations and becomes computationally prohibitive at the fine voxel resolution employed by DCA. Therefore, we computed DCBC scores by projecting volumetric atlases onto the cortical surface (fsLR 32k template [53]), analyzing only data from the left hemisphere. This surface-based approach exceeds the scope of our native volumetric framework (Fig. 9 and Table 11 ).
@@
Atlas #Parcels mean std.
@@ 14 Effect of parcel number on downstream tasks
To examine whether downstream tasks prefer specific spatial resolutions, we varied the number of parcels for both DCA (41, 100, 200, 360, 400, 500) and Schaefer (100, 200, 300, 400, 500) atlases. For each resolution, we performed subject-level cross-validation and report accuracies in Tables 13 and 14. Across both atlas families, three consistent patterns emerge:
@@
  • Resolution-insensitive tasks. Several clinical endpoints (e.g., AD diagnosis) exhibit weak or no monotonic trend across resolutions for both DCA and Schaefer. These tasks likely rely on subcortical or global signals that cortex-only atlases do not explicitly model, so varying cortical granularity alone has a limited effect. In such settings, choosing resolution can be guided by computational cost or downstream interpretability rather than accuracy.
@@ 15.1 Regularization and reconstruction loss
We investigated whether incorporating an orthogonality regularizer or a reconstruction loss would further improve parcellation quality (Table 15). Applying these augmented objectives to the same 100 subjects used in our main experiments yielded no statistically significant gains in homogeneity or silhouette coefficient for 100 parcels, indicating that the core KL‐based clustering loss is sufficient to drive optimal voxel‐level atlas generation.
@@
The orthogonality regularizer is defined on the centroid matrix $\mathbf{D}\in\mathbb{R}^{K\times d}$ (with unit‐norm rows) as follows. Let
@@
be the Gram matrix of pairwise inner products. We zero out the diagonal to isolate off‐diagonal similarities:
@@
The orthogonality loss then penalizes the mean absolute off‐diagonal entry via
@@ 15.3 Choice of graph cut method
In addition to spectral clustering, we evaluated several graph‐cut algorithms (Fig. 10 and Table 17). However, most failed to guarantee that each resulting parcel forms a single connected subgraph, leading to fragmented regions. Here, we use a breadth-first search (BFS) based algorithm. The weighted BFS–connected clustering algorithm begins by converting the input edge list and weights into an undirected adjacency list, then randomly seeds $k$ initial clusters by assigning one unique node to each cluster. Each cluster maintains a max‐heap of its unassigned neighboring nodes, prioritized by edge weight. Clusters then grow in parallel: at each step, a cluster pops the highest‐weight neighbor from its heap, claims that node (if unassigned), and pushes all of its unassigned neighbors onto the heap. To enforce roughly equal cluster sizes, each cluster stops growing once it reaches $\lceil N/k\rceil$. If any nodes remain unassigned after this frontier‐driven expansion, they are absorbed into the smallest adjacent cluster. By always selecting the strongest edges first and only adding connected nodes, this method produces contiguous clusters that respect the underlying graph’s weighted connectivity.